For the reaction

N₂O₄(g) ⇔ 2NO₂(g)

K_P = 1.4×10⁴Pa at 303K. To find the value of K_c at this temperature.


1
Expert's answer
2021-02-15T03:00:27-0500

From the relationship,

"Kp = Kc(RT)^n"

1.4×104 = Kc(0.0821×303)

Kc = [563] M